# Vision-Based Mistake Analysis in Procedural Activities: A Review of Advances and Challenges

### v1

@@ IV-A Datasets: overview & key attributes
EgoOops [12]: The dataset consists of 40 video recordings from 5 diverse task domains, namely electrical circuit assembly, color mixture experiments, ionic reactions, toy block construction and cardboard crafts, captured in authentic but constrained (in terms of environment), educational and workshop settings. A key feature of EgoOops, shared only by CaptainCook4D and EgoPER, is procedural text integration, with texts tightly aligned with video segments to facilitate step-level grounding. EgoOops offers segment-wise annotations, including video-text alignments, detailed mistake labels (covering order and execution mistakes), object label annotations (microQR) and natural language descriptions of why specific actions deviate from the intended procedure, allowing for mistake explainability. Procedural mistakes in EgoOops follow the common ordering error scenarios found in related datasets, such as skipped, swapped, or repeated steps. In addition, the dataset introduces a detailed taxonomy of 6 executional mistakes, which occur when participants fail to correctly follow instructions in terms of executing the steps of an action (e.g., grasp wrong object). Finally, participants were instructed to follow scripted task and error scenarios, ensuring controlled coverage of error types; however, unintentional errors also occurred naturally during execution and were also annotated.
@@ IV-B Datasets: Comparison & Discussion
EpicTent [17] 2018, 2023 Assembly/Real Ego 1 38 MR PEs∗∗ NA 24 >>5.4h 29 RGB, Gaze
CSV [18] 2022 Chemistry/Lab Ego 14 106 MR, SR, EMR PEs NA 70 11.1h 82 RGB
Assembly101 [14] 2022 Toy Assembly/Lab Ego, Exo 101 202 MR, ASD, SR PEs, EEs 15 362 167.0h 53 RGB, 3D Hand Pose
BRIO-TA [19] 2022 Assembly Exo 1 23 MR, SR PEs 10 75 2.9h 15 RGB
ATA [20] 2023 Toy Assembly/Lab Exo 3 15 MR, SR PEs 11 141 24.8h 32 RGB, Audio, Text, Pose, Obj. BB
HoloAssist [21] 2023 Assembly/Lab Ego 20 414 MR, SR, ASD, SA, EMR PEs, EEs 16 350 166h 222 RGB, Depth, Pose (Hand & Head), Gaze, Text, IMU
CaptainCook4D [13] 2023 Cooking/Real Ego 24 352 MR, SR, EMR, TAML PEs, EEs NA 384 94.5h 8 RGB, Depth, Text, Task Graphs
Ego-Exo4D[25] 2024 Physical Exercise, Assembly/Real Exo, Ego 6 186 MR, SR, SA PEs 5.6K∗∗∗ 628 30h 740 RGB, Audio, Gaze, Depth, IMU, Text, Task Graphs, Bar, Mag
IndustReal[23] 2024 Toy Assembly/Lab Ego 2 75 MR, ASD, SR, TAML PEs, EEs 36 84 5.8h 27 RGB, Depth, Gaze, Hand, Pose
EgoPER[15] 2024 Cooking Ego 5 70 MR, TAML PEs, EEs 35 386 28h 11 RGB, Depth, Audio, Gaze, Hand BB, Obj BB, Task Graphs
EgoOops[12] 2024 Diverse Ego 5 46 MR, TAML PEs, EEs 58 40 6.8h 4 RGB, Text, Object (label-only)
TABLE I: Procedural understanding datasets targeting mistake detection. Supported tasks include MR: mistake recognition, ASD: activity state detection, SR: step (action) recognition, TAML: temporal action and mistake localization, SA: skill assessment, EMR: early mistake recognition. Partps: Participants. Text refers to transcript availability. Additional abbreviations: PEs: procedural errors, $E$s: execution errors, 3DM: 3D models publicly available, OF: Optical Flow, $B$: bounding boxes, Bar: Barometer, Mag: Magnetometer. sparse annotations, ∗∗ refined in subsequent works, ∗∗∗ refers to total objects annotated across all dataset-supported tasks, only a subset used for mistake analysis (exact number not provided), NA: not annotated.
@@
Exceptions are the variant of Assembly101 proposed by Ding et al. [16] and, more recently, the IndustReal dataset. The first incorporates the notion of error accumulation in the mistake detection task using a rule-based action transition analysis scheme, where an action transition rule is classified as transitive or intransitive, depending on how mistakes propagate through the action sequence. Transitive rules imply that any subsequent dependent action after an incorrect anchor action is also a mistake, while intransitive rules consider only specific dependent actions as mistakes. Although very generalizable, this approach is accompanied by physical annotations of specific cases in the dataset. Contrary, IndustReal emphasizes long-horizon task monitoring in real industrial environments, where an initial error can propagate through subsequent actions. Its design supports tracing such error chains, offering a realistic depiction of procedural degradation over time, crucial for systems aiming to assist real-world task execution.
@@ V-A1 Mistake Recognition
The predominant evaluation strategy adopted in prior work [15] involves formulating mistake recognition either as a binary classification task, wherein each action is labeled as correct or mistaken, or as a multi-class classification task that further distinguishes between specific error types. In these settings, standard classification metrics such as precision, recall, F1 score and the area under the receiver operating characteristic curve (AUC) are commonly used to quantify model performance.
While these general-purpose metrics provide a foundational assessment, several works have proposed adaptations that better reflect the temporal and procedural nature of the mistake detection task. One such task-specific metric is the Error Detection Accuracy (EDA) [15, 26], originally introduced in [15]. EDA is defined as the ratio of correctly detected erroneous segments over the total number of ground-truth erroneous segments across all test videos. segments. Specifically, given the test set, $V_{t}$, consisting of $K$ videos, $V_{t}=\{V_{1},\dots,V_{K}\}$, where each video $V_{k}$ is decomposed into a set of segments $V_{k}=\{v_{i}^{(k)}\}_{i=1}^{N_{k}}$, with corresponding ground-truth mistake labels $m_{i}^{(k)}\in\{0,1\}$ and predicted mistake labels $\hat{m}_{i}^{(k)}\in\{0,1\}$, the EDA is defined as:
@@
A related metric, the Anomaly Segment Accuracy (ASA), was introduced in the BRIO-TA dataset [19] to quantify segment-level performance in mistake recognition and detection tasks. ASA is a segment-level analogue of the standard accuracy metric, adapted to mistake analysis. The ASA metric measures the proportion of action segments that are correctly classified as either normal or erroneous, and is defined as
@@ V-A2 Early Mistake Recognition
This task resembles early action recognition, where the goal is to anticipate an event before it fully unfolds. Therefore, similar protocols and evaluation metrics have been partially adopted. In early action recognition [27], common strategies include measuring classification accuracy at fixed observation percentages (e.g., $10\%$, $20\%$, …, $100\%$ of the action duration, termed anticipation windows) or computing metrics like AUC to capture performance over time. Similarly, recent works in early mistake recognition [13], [18], [21] evaluated model performance at different anticipation windows, gauging predictive accuracy before the full mistake occurred. As an example, CaptainCook4D[13] introduces an early error detection task, where models are evaluated at fixed intervals prior to the annotated mistake point (e.g., 2s before). Performance is measured using Top-1 accuracy and F1-score within these anticipation windows. This allows for assessing how well the model can detect an impending mistake with limited temporal context. In a related but slightly different fashion, SVIP [28] evaluates sequence verification models on partially observed video snippets to assess the ability to recognize incomplete or incorrect procedures early on. While SVIP focuses on binary classification accuracy of a completed vs. incomplete (or incorrect) procedure, its use of early segment evaluation reflects similar motivations to those found in early action recognition literature.
@@ V-A3 Mistake Detection
As outlined in Section VI, mistake detection can be formulated as a temporal action localization (TAL) problem, where the objective is to identify the type of mistake and its temporal boundaries within the video. Under this format, several recent works [12], [13, 19] adopt evaluation protocols derived from established TAL benchmarks [29, 30], employing metrics such as mean Average Precision (mAP) and Recall at X (R@X). These metrics are computed at varying temporal Intersection-over-Union (tIoU) thresholds, typically ranging from 0.1 to 0.5 in increments of 0.1 or 0.2.
@@
Beyond conventional TAL metrics, some recent works have introduced task-specific measures that capture higher-order aspects of procedural consistency. Notably, Schoonbeek et al. [23] propose the Procedure Order Similarity (POS) metric to assess how well the predicted sequence of completed steps adheres to the canonical step order. POS is defined as:
@@
where $\mathcal{P}=\{a_{i}\}_{i=1}^{N}$ is the ordered ground-truth step sequence, $\hat{\mathcal{P}}=\{\hat{a}_{j}\}_{j=1}^{M}$ is the predicted sequence, and DamLev denotes the Damerau–Levenshtein distance [31] without substitutions. This metric accounts for missing, repeated, or wrongly ordered steps providing a holistic measure of procedural correctness.
@@ V-B Mistake/Error Type-specific Metrics
Omission errors using action frequencies: The work of Ghoddoosian et al. [20] proposed a metric to evaluate the performance of mistake detection methods by expressing the discrepancies between expected and predicted action frequencies for specific action pairs. The metric relies on defining error-specific functions $\{F_{e}\}$ that operate over these frequencies. Each function maps the observed frequency $f_{a}$ of actions in the test video to the number of instances $n_{e}$ in which an error $e$ occurs. For example, the error Loose Assembly, arising when a component is positioned but not secured, is expressed as:
@@
Omission errors using edit distance: Lee et al. [15] introduced the Omission Intersection over Union (O-IoU) metric to evaluate the performance of mistake detection methods, particularly in scenarios involving omissions, defined as:
@@
Procedural error assessment using Weighted Distance Ratio (WDR): Qian et al. [18] introduced WDR as an evaluation metric for procedural mistake detection, which quantifies the performance of a method by assessing the similarity between action sequences from correct and incorrect procedural executions. WDR is defined as the ratio of the average distance between negative video pairs to the average distance between positive video pairs. A positive pair consists of two videos in which the procedural steps are executed correctly and in the correct order, whereas a negative pair contains at least one video with a procedural error. The metric is defined as:
@@ VI-A Mistake Detection and Recognition
Several works [13, 12, 20] tackled mistake detection as a temporal action localization problem, assuming that deviations from expected procedural flow, such as missing steps, incorrect orderings, or executional anomalies, manifest as misaligned or inconsistent temporal patterns. Most approaches build upon self- or fully supervised action segmentation models fine-tuned on task-specific action pools to segment untrimmed videos into procedural steps. Mistake presence is then inferred through misalignment with expected sequences, unexpected transitions, or out-of-distribution behavior. Methods such as [13, 15] employ spatiotemporal backbones (I3D [32], SlowFast [33]) with segmentation or detection modules (e.g., ActionFormer [34], MS-TCN++ [35], MiniRoad [36]) followed by mistake recognition processes like segment-level classification [13], misalignment scoring [15], or temporal inconsistency detection [12]. Some also exploit text-based procedural priors alongside RGB to identify temporal inconsistencies—for example, EgoOops [12] aligns predicted actions with task descriptions to highlight erroneous segments. In weakly supervised settings [20], models rely solely on ordered transcripts and constrained alignment objectives to localize unseen mistakes without direct annotations. Despite architectural differences, these methods share a common goal: to detect when mistakes occur in untrimmed video and to segment the corresponding intervals.
Recent works reformulate this offline setting into an online pipeline, making predictions incrementally as new frames arrive, using only past context. Flaborea et al. [24] introduced PREGO, combining an online action recognition module with a symbolic reasoning branch driven by a pre-trained LLM that predicts the next expected action and flags discrepancies as mistakes. Plini et al. [37] extend this by adding chain-of-thought reasoning and few-shot in-context examples, improving adaptability and temporal precision.

### v2

@@ IV-A Datasets: overview & key attributes
EgoOops [EgoOops]: The dataset consists of 40 video recordings from 5 diverse task domains, namely electrical circuit assembly, color mixture experiments, ionic reactions, toy block construction and cardboard crafts, captured in authentic but constrained (in terms of environment), educational and workshop settings. A key feature of EgoOops, shared only by CaptainCook4D and EgoPER, is procedural text integration, with texts tightly aligned with video segments to facilitate step-level grounding. EgoOops offers segment-wise annotations, including video-text alignments, detailed mistake labels (covering order and execution mistakes), object label annotations (microQR) and natural language descriptions of why specific actions deviate from the intended procedure, allowing for mistake explainability. Procedural mistakes in EgoOops follow the common ordering error scenarios found in related datasets, such as skipped, swapped, or repeated steps. In addition, the dataset introduces a detailed taxonomy of 6 executional mistakes, which occur when participants fail to correctly follow instructions in terms of executing the steps of an action (e.g., grasp wrong object). Finally, participants were instructed to follow scripted task and error scenarios, ensuring controlled coverage of error types; however, unintentional errors also occurred naturally during execution and were also annotated.
Ego4D-M & EPIC-KITCHENS-M: Addressing the scarcity of large-scale mistake data, Li et al. [li2025mistake] introduced EK-M and Ego4D-M, which were synthetically created by repurposing existing egocentric corpora via a data engine named MisEngine. Unlike datasets relying on staged or naturally occurring errors, these benchmarks are generated by systematically cross-matching instruction texts with video segments of disparate actions (e.g., pairing a “pick up hammer” instruction with a “pick up bolt” video) based on Semantic Role Labeling (SRL) mismatches. This automated misalignment process allows for the inheritance of rich annotations from the source datasets (Ego4D [grauman2022ego4d] and EPIC-KITCHENS(EK) [damen2022rescaling]), resulting in benchmarks that are two orders of magnitude larger than prior works. Specifically, they provide approximately 257K and 221K samples respectively, covering $12,283$ and $16,099$ action classe respectively, offering supervision for fine-grained semantic, temporal (frame-level PNR  [^1] Point-of-No -Return: signifies the reference frame for the mistake.), and spatial (bounding box) mistake attribution.
@@ IV-B Datasets: Comparison & Discussion
EpicTent [EPICTENT] 2018, 2023 Assembly/Real Ego 1 38 MR PEs∗∗ NA 24 >>5.4h 29 RGB, Gaze
CSV [CSV] 2022 Chemistry/Lab Ego 14 106 MR, SR, EMR PEs NA 70 11.1h 82 RGB
Assembly101 [Assembly101] 2022 Toy Assembly/Lab Ego, Exo 101 202 MR, ASD, SR PEs, EEs 15 362 167.0h 53 RGB, 3D Hand Pose
BRIO-TA [BRIOTA] 2022 Assembly Exo 1 23 MR, SR PEs 10 75 2.9h 15 RGB
ATA [ATA] 2023 Toy Assembly/Lab Exo 3 15 MR, SR PEs 11 141 24.8h 32 RGB, Audio, Text, Pose, Obj. BB
HoloAssist [HoloAssist] 2023 Assembly/Lab Ego 20 414 MR, SR, ASD, SA, EMR PEs, EEs 16 350 166h 222 RGB, Depth, Pose (Hand & Head), Gaze, Text, IMU
CaptainCook4D [CaptainCook4D] 2023 Cooking/Real Ego 24 352 MR, SR, EMR, TAML PEs, EEs NA 384 94.5h 8 RGB, Depth, Text, Task Graphs
Ego-Exo4D[grauman2024ego] 2024 Physical Exercise, Assembly/Real Exo, Ego 6 186 MR, SR, SA PEs 5.6K∗∗∗ 628 30h 740 RGB, Audio, Gaze, Depth, IMU, Text, Task Graphs, Bar, Mag
IndustReal[IndustReal] 2024 Toy Assembly/Lab Ego 2 75 MR, ASD, SR, TAML PEs, EEs 36 84 5.8h 27 RGB, Depth, Gaze, Hand, Pose
EgoPER[EgoPED] 2024 Cooking/Real Ego 5 70 MR, TAML PEs, EEs 35 386 28h 11 RGB, Depth, Audio, Gaze, Hand BB, Obj BB, Task Graphs
EgoOops[EgoOops] 2024 Diverse/Real Ego 5 46 MR, TAML PEs, EEs 58 40 6.8h 4 RGB, Text, Object (label-only)
EK-M[li2025mistake] 2025 Cooking/Real(synth.) Ego Inherited 12.3K MR, TAML PEs, EEs Inherited >>221K NATW 37 RGB, Text
Ego4D-M[li2025mistake] 2025 Diverse/Real(synth.) Ego Inherited 16K MR, STAML PEs, EEs Inherited >>257K NATW 248 RGB, Text, PNR, BB
TABLE I: Procedural understanding datasets targeting mistake detection. Supported tasks: MR: mistake recognition, ASD: activity state detection, SR: step (action) recognition, TAML: temporal action and mistake localization, STAML: spatiotemporal action and mistake localization, SA: skill assessment, EMR: early mistake recognition. Partps: Participants. Text refers to transcript availability. Note that: TAML and STAML correspond to the general task of Mistake Detection (MD). Additional abbreviations: PEs: procedural errors, $E$s: execution errors, 3DM: 3D models publicly available, OF: Optical Flow, $B$: bounding boxes, Bar: Barometer, Mag: Magnetometer. sparse annotations, ∗∗ refined in subsequent works, ∗∗∗ refers to total objects annotated across all dataset-supported tasks, only a subset used for mistake analysis (exact number not provided), NA: not annotated, NATW: non-available at the time of writing. Inherited: the synthetic dataset inherits the specific attribute from the original.
@@
Exceptions are the variant of Assembly101 proposed by Ding et al. [ding2023mistakecountsassembly] and, more recently, the IndustReal dataset. The first incorporates the notion of error accumulation in the mistake detection task using a rule-based action transition analysis scheme, where an action transition rule is classified as transitive or intransitive, depending on how mistakes propagate through the action sequence. Transitive rules imply that any subsequent dependent action after an incorrect anchor action is also a mistake, while intransitive rules consider only specific dependent actions as mistakes. Although very generalizable, this approach is accompanied by physical annotations of specific cases in the dataset. Contrary, IndustReal emphasizes long-horizon task monitoring in real industrial environments, where an initial error can propagate through subsequent actions. Its design supports tracing such error chains, offering a realistic depiction of procedural degradation over time, crucial for systems aiming to assist real-world task execution.
@@ V-A1 Mistake Recognition
The predominant evaluation strategy adopted in prior work [EgoPED] involves formulating mistake recognition either as a binary classification task, wherein each action is labeled as correct or mistaken, or as a multi-class classification task that further distinguishes between specific error types. In these settings, standard classification metrics such as precision, recall, F1 score and the area under the receiver operating characteristic curve (AUC) are commonly used to quantify model performance.
While these general-purpose metrics provide a foundational assessment, several works have proposed adaptations that better reflect the temporal and procedural nature of the mistake detection task. One such task-specific metric is the Error Detection Accuracy (EDA) [EgoPED, kung2025changed], originally introduced in [EgoPED]. EDA is defined as the ratio of correctly detected erroneous segments over the total number of ground-truth erroneous segments across all test videos. segments. Specifically, given the test set, $V_{t}$, consisting of $K$ videos, $V_{t}=\{V_{1},\dots,V_{K}\}$, where each video $V_{k}$ is decomposed into a set of segments $V_{k}=\{v_{i}^{(k)}\}_{i=1}^{N_{k}}$, with corresponding ground-truth mistake labels $m_{i}^{(k)}\in\{0,1\}$ and predicted mistake labels $\hat{m}_{i}^{(k)}\in\{0,1\}$, the EDA is defined as:
@@
A related metric, the Anomaly Segment Accuracy (ASA), was introduced in the BRIO-TA dataset [BRIOTA] to quantify segment-level performance in mistake recognition and detection tasks. ASA is a segment-level analogue of the standard accuracy metric, adapted to mistake analysis. The ASA metric measures the proportion of action segments that are correctly classified as either normal or erroneous, and is defined as
@@ V-A2 Early Mistake Recognition
This task resembles early action recognition, where the goal is to anticipate an event before it fully unfolds. Therefore, similar protocols and evaluation metrics have been partially adopted. In early action recognition [zhao2021review], common strategies include measuring classification accuracy at fixed observation percentages (e.g., $10\%$, $20\%$, …, $100\%$ of the action duration, termed anticipation windows) or computing metrics like AUC to capture performance over time. Similarly, recent works in early mistake recognition [CaptainCook4D], [CSV], [HoloAssist] evaluated model performance at different anticipation windows, gauging predictive accuracy before the full mistake occurred. As an example, CaptainCook4D[CaptainCook4D] introduces an early error detection task, where models are evaluated at fixed intervals prior to the annotated mistake point (e.g., 2s before). Performance is measured using Top-1 accuracy and F1-score within these anticipation windows. This allows for assessing how well the model can detect an impending mistake with limited temporal context. In a related but slightly different fashion, SVIP [SVIP] evaluates sequence verification models on partially observed video snippets to assess the ability to recognize incomplete or incorrect procedures early on. While SVIP focuses on binary classification accuracy of a completed vs. incomplete (or incorrect) procedure, its use of early segment evaluation reflects similar motivations to those found in early action recognition literature.
@@ V-A3 Mistake Detection
As outlined in Section VI, mistake detection can be formulated as a temporal action localization (TAL) problem, where the objective is to identify the type of mistake and its temporal boundaries within the video. Under this format, several recent works [EgoOops], [CaptainCook4D, BRIOTA] adopt evaluation protocols derived from established TAL benchmarks [lai2024human, zhong2023survey], employing metrics such as mean Average Precision (mAP) and Recall at X (R@X). These metrics are computed at varying temporal Intersection-over-Union (tIoU) thresholds, typically ranging from 0.1 to 0.5 in increments of 0.1 or 0.2.
@@
Beyond conventional TAL metrics, some recent works have introduced task-specific measures that capture higher-order aspects of procedural consistency. Notably, Schoonbeek et al. [IndustReal] propose the Procedure Order Similarity (POS) metric to assess how well the predicted sequence of completed steps adheres to the canonical step order. POS is defined as:
@@
where $\mathcal{P}=\{a_{i}\}_{i=1}^{N}$ is the ordered ground-truth step sequence, $\hat{\mathcal{P}}=\{\hat{a}_{j}\}_{j=1}^{M}$ is the predicted sequence, and DamLev denotes the Damerau–Levenshtein distance [damerau1964technique] without substitutions. This metric accounts for missing, repeated, or wrongly ordered steps providing a holistic measure of procedural correctness.
@@ V-B Mistake/Error Type-specific Metrics
Omission errors using action frequencies: The work of Ghoddoosian et al. [ATA] proposed a metric to evaluate the performance of mistake detection methods by expressing the discrepancies between expected and predicted action frequencies for specific action pairs. The metric relies on defining error-specific functions $\{F_{e}\}$ that operate over these frequencies. Each function maps the observed frequency $f_{a}$ of actions in the test video to the number of instances $n_{e}$ in which an error $e$ occurs. For example, the error Loose Assembly, arising when a component is positioned but not secured, is expressed as:
@@
Omission errors using edit distance: Lee et al. [EgoPED] introduced the Omission Intersection over Union (O-IoU) metric to evaluate the performance of mistake detection methods, particularly in scenarios involving omissions, defined as:
@@
Procedural error assessment using Weighted Distance Ratio (WDR): Qian et al. [CSV] introduced WDR as an evaluation metric for procedural mistake detection, which quantifies the performance of a method by assessing the similarity between action sequences from correct and incorrect procedural executions. WDR is defined as the ratio of the average distance between negative video pairs to the average distance between positive video pairs. A positive pair consists of two videos in which the procedural steps are executed correctly and in the correct order, whereas a negative pair contains at least one video with a procedural error. The metric is defined as:
@@
V-C Mistake Explanation Metrics
With the emergence of methods that provide natural language justifications for detected errors [patsch2025mistsense], evaluating the semantic quality and diagnostic accuracy of these explanations has become a critical sub-task. Current approaches primarily repurpose standard image and video captioning metrics to assess the similarity between the generated explanation and a ground-truth reference.
The most common evaluation protocols utilize N-gram overlap metrics, including BLEU[papineni2002bleu], ROUGE[lin-2004-rouge], and CIDEr[vedantam2015cider]. These metrics quantify performance by calculating the precision and recall of word sequences between the predicted and reference explanations. For instance, Patsch et al.[patsch2025mistsense] employ CIDEr as a primary metric to capture the consensus of generated explanations with human annotations. While these metrics effectively measure linguistic fluency and textual similarity, they can fail to capture diagnostic correctness in the context of mistake analysis. A generated explanation could achieve a high BLEU score by matching the majority of a sentence (e.g., “The user failed to pick up the object”) while missing or hallucinating the critical causal detail (e.g., specifying the wrong object).
@@ VI-A Mistake Detection and Recognition
Several works [CaptainCook4D, EgoOops, ATA] tackled mistake detection as a temporal action localization problem, assuming that deviations from expected procedural flow, such as missing steps, incorrect orderings, or executional anomalies, manifest as misaligned or inconsistent temporal patterns. Most approaches build upon self- or fully supervised action segmentation models fine-tuned on task-specific action pools to segment untrimmed videos into procedural steps. Mistake presence is then inferred through misalignment with expected sequences, unexpected transitions, or out-of-distribution behavior. Methods such as [CaptainCook4D, EgoPED] employ spatiotemporal backbones (I3D [carreira2017quo], SlowFast [feichtenhofer2019slowfast]) with segmentation or detection modules (e.g., ActionFormer [zhang2022actionformer], MS-TCN++ [farha2019ms], MiniRoad [an2023miniroad]) followed by mistake recognition processes like segment-level classification [CaptainCook4D], misalignment scoring [EgoPED], or temporal inconsistency detection [EgoOops]. Some also exploit text-based procedural priors alongside RGB to identify temporal inconsistencies, for example, EgoOops [EgoOops] aligns predicted actions with task descriptions to highlight erroneous segments. In weakly supervised settings [ATA], models rely solely on ordered transcripts and constrained alignment objectives to localize unseen mistakes without direct annotations. Despite architectural differences, these methods share a common goal: to detect when mistakes occur in untrimmed video and to segment the corresponding intervals.
A recent work by Li et al. [li2025mistake] pushes the boundaries of offline mistake detection beyond simple temporal intervals by introducing Mistake Attribution (MATT). Unlike standard detection schemes followed by existing methods, which flag the entire duration of a deviation (entire action segment (procedural error) or a portion of the action segment (executional error)), MATT aims to spatiotemporally localize the mistake occurrence. To achieve this, they introduce a mechanism to detect the Point-of-No-Return (PNR) frame, which corresponds to the exact instant the mistake becomes irreversible. The PNR serves as a temporal anchor for spatial reasoning, enabling models to localize the mistake’s visual manifestation (e.g., via bounding boxes) on the exact frame where the error consolidates.
Finally, recent works attempt to reformulate this offline setting into an online pipeline, making predictions incrementally as new frames arrive, using only past context. Flaborea et al. [flaborea2024prego] introduced PREGO, combining an online action recognition module with a symbolic reasoning branch driven by a pre-trained LLM that predicts the next expected action and flags discrepancies as mistakes. Plini et al. [plini2024ti] extend this by adding chain-of-thought reasoning and few-shot in-context examples, improving adaptability and temporal precision.